1). Determine whether set S given below is a basis for ℝ 3 . If not, explain why.

S = {(1,0,0),(0,1,0),(0,0,0)}


2). Find the rank of matrix D given below.

D = [0 3 9 0]

[-2 -1 1 -1]

[0 -1 -3 1]


1
Expert's answer
2021-09-16T00:43:30-0400

a) A subset "S" of a vector space "V" is called a basis if

1. "S" is linearly independent, and

2. "S"  is a spanning set

Any three linearly independent vectors form a basis of "\R^3."

Let us check that whether "S" is a linearly independent set.

Consider the linear combination


"x_1\begin{pmatrix}
 1 \\
 0 \\
 0
\end{pmatrix}+x_2\begin{pmatrix}
 0 \\
 1 \\
 0
\end{pmatrix}+x_3\begin{pmatrix}
 0 \\
 0 \\
 0
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
 0
\end{pmatrix}"

This is equivalent to the matrix equation


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 0 \\
\end{pmatrix}\begin{pmatrix}
 x_1 \\
 x_2 \\
 x_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
 0
\end{pmatrix}""A=\begin{pmatrix}
 1 & 0 & 0 & & 0 \\
 0 & 1 & 0 & & 0 \\
 0 & 0 & 0 & & 0 \\
\end{pmatrix}"

Thus, the general solution is "x_1=0, x_2=0, x_3=t, t\in \R" is a free variable.

Hence, in particular, there is a nonzero solution.

So "S" is linearly dependent, and hence "S" cannot be a basis for "\R^3."


b)


"D= \begin{bmatrix}
 0 & 3 & 9 & 0 \\
 -2 & -1 & 1 & -1 \\
 0 & -1 & -3 & 1 \\
\end{bmatrix}"

Swap "R_1" and "R_2"


"\begin{bmatrix}
 -2 & -1 & 1 & -1 \\
 0 & 3 & 9 & 0 \\
 0 & -1 & -3 & 1 \\
\end{bmatrix}"

"R_3=R_3+R_2\/3"


"\begin{bmatrix}
 -2 & -1 & 1 & -1 \\
 0 & 3 & 9 & 0 \\
 0 & 0 & 0& 1 \\
\end{bmatrix}"

The rank of a matrix is the number of nonzero rows in the reduced matrix, so the rank is 3.


"rank(D)=3"